OpenStudy (anonymous):

What is the slope of a line parallel to the line whose equation is 2x – 3y = 6? A. 3/3 B. 4/3 C. 2/3 D. 1/5

OpenStudy (anonymous):

a

OpenStudy (anonymous):

k thx @hatemath453

OpenStudy (anonymous):

hey can u help wit another 1

OpenStudy (noorik.):

wrong it's c

OpenStudy (anonymous):

k

OpenStudy (noorik.):

because the linear equation is y=mx+b m= slope 2x-3y=6 2x-6=3y 2x/3 - 2 = y 2/3=m